# Blowing up Dirac’s theorem

### v1

@@ 6. Connecting and rooting blow-ups
In this section, we show Lemmata 2.2 and 4.2, starting with the former.
@@
Note that every vertex in $U\cup V$ has at least $(1/2+\varepsilon/2)n$ neighbours in $W$. Let $W_{1}\subseteq W$ be the set of vertices of $W$ with at least $\varepsilon m/8$ neighbours in $U$. Counting edges between $U$ and $W$, we get $(1/2+\varepsilon/2)mn\leq|W_{U}|m+(n-|W_{U}|)\varepsilon m/8$, which gives
@@
Now let $W_{V}\subseteq W$ be the set of vertices of $W$ with at least $\varepsilon m/8$ neighbours in $V$. An analogous argument gives that $|W_{U}|\geq\left(1/2+\varepsilon/2\right)n$. Hence, the set $W^{\ast}=W_{U}\cap W_{V}$ has size at least $\varepsilon n$.
Set $t=m^{\prime}$. We claim that there are $t$-sets $W_{1}\subseteq U$, $W_{2}\subseteq V$ and $W_{3}\subseteq W$ such that $G$ contains the complete bipartite graphs $K(W_{1},W_{3})$ and $K(W_{2},W_{3})$. Suppose otherwise. So no two $t$-sets in $U$ and $V$ have more than $t-1$ common neighbours in $W^{\ast}$. Therefore
@@
It remains to show how to obtain rooted blow-ups.
@@
Let $G^{\prime}$ be the graph obtained from $G$ by removing the edges inside $V$. Since $|V|=m_{1}\leq\varepsilon n/4$ this barely affects the minimum degree. So we still have $\delta(G^{\prime})\geq(1/2+3\varepsilon/4)n$. Let $P$ be the $s$-uniform hypergraph where an $s$-set forms an edge if it has exactly one vertex in $V$ and the induced subgraph $G^{\prime}[S]$ has minimum degree at least $(1/2+\varepsilon/2)s$. By Lemma 3.1 and crude bounding, we have $e(P)\geq|V|n^{s-1}/2$. Consider a random partition of $V(G)\smallsetminus V$ by assigning each vertex uniformly to one of $s-1$ parts. Together with $V$, this gives a partition of $V(G)$ into $s$ parts. Let $P^{\prime}\subseteq P$ be the $s$-partite $s$-graph, which contains the crossing edges. By a first moment argument, it follows that with positive probability $e(P^{\prime})\geq|V|n^{s-1}k!/(2k^{k})$. Fix such a partition for the rest of the proof. Recall that there are at most $2^{s^{2}}$ labelled graphs on $s$ vertices. Let $R$ be the labelled graph that appears on the edges of $P^{\prime}$ most frequently, where the labels correspond to the parts. We pass to a subgraph $P^{\prime\prime}\subseteq P^{\prime}$ keeping only the $s$-sets $S$ where $G[S]$ is isomorphic to $R$. So $P^{\prime\prime}$ has at least $|V|n^{s-1}k!/(k^{k}2^{s^{2}+2})$ edges. Note that by our choice we can assume that for each $S\in P^{\prime\prime}$, $G[S]$ is isomorphic to $R$, $|S\cap V|=1$ and also $G[S\smallsetminus V]$ is isomorphic to the same $(s-1)$-graph $R^{\prime}$, no matter the choice of $S$.
@@
Now let $1/s\gg\gamma_{1}\gg\gamma_{2}\gg c$, so that $e(P^{\prime\prime})\geq\gamma_{1}|V|n^{s-1}$. Let $\mathcal{R}^{\prime}$ be the set of ordered $(s-1)$-tuples in $G-V$, inducing a copy of $R^{\prime}$, which can be extended to at least $\gamma_{1}|V|/2$ copies of $R$ by adding a vertex of $V$. Since $e(P^{\prime\prime})\geq\gamma_{1}|V|n^{s-1}$, we have that $|\mathcal{R}^{\prime}|\geq\gamma_{1}n^{s-1}/2$. By Lemma 3.2, there is a blow-up $K\subseteq G-V$ of $R^{\prime}$ with all clusters of size $t=\gamma_{2}\log n$. Let $R^{\prime}_{1},\dotsc,R^{\prime}_{t}$ be vertex-disjoint copies of $R^{\prime}$ in the blow-up $K$. Form an auxiliary bipartite graph $B$ with vertex classes $V$ and $W=\{R^{\prime}_{1},\dotsc,R^{\prime}_{t}\}$ where $u\sim R^{\prime}_{i}$ if $R^{\prime}_{i}$ can be extended to a copy of $R$ by adding $u$. Then $B$ has at least $\gamma_{1}t|V|/2$ edges.
@@
Note that to conclude it is enough to find a complete bipartite graph in $B$ with parts of size $c\ln n$. The existence of such graph follows from the usual bounds for the Zarankiewicz problem. To be precise, let $p=c\log n$ and suppose $B$ does not have a complete bipartite graph with two classes of size $p$. Then each $p$-set $X\subseteq W$ has at most $p-1$ common neighbours $N_{B}(X)$ in $V$. Thus, double-counting and convexity reveals
$(p-1)\binom{t}{p}$ $\geq\sum_{X\in\binom{W}{p}}|N_{B}(X)|=\sum_{v\in V}\binom{\deg_{B}(v)}{p}$
$\geq|V|\binom{e(B)/|V|}{p}\geq|V|\binom{\gamma_{1}t/2}{p}\geq(\gamma_{1}/4)^{p}|V|\binom{t}{p},$
@@
$p\geq(\gamma_{1}/4)^{p}|V|\geq|V|n^{-1/4}\geq n^{1/4},$
where the second inequality follows from the (small enough) choice of $c$, and the last inequality follows from $|V|=m_{1}\geq n^{1/2}$. This contradicts the choice of $p$, and proves the result. ∎
@@ Acknowledgements
Richard Lang was supported by the Ramón y Cajal programme (RYC2022-038372-I) and by grant PID2023-147202NB-I00 funded by MICIU/AEI/10.13039/501100011033. Nicolás Sanhueza-Matamala was supported by ANID-FONDECYT Iniciación Nº11220269 grant.
@@ A. Inheritance
For a $k$-graph $G$ on $n$ vertices, we denote the relative degree of vertex $v\in V(G)$ by $\overline{\deg}_{G}(v)=\deg_{G}(v)/(n-1)$. The following instance of the inheritance principle immediately implies Lemma 3.1.
@@
We show Lemma A.1 using the following standard tail bound. Recall that a hypergeometrically distributed random variable $X\sim\text{H}(N,K,n)$ describes the probability of $k$ successes in $n$ draws, without replacement, from a finite population of size $N$ that contains exactly $K$ objects with that feature, wherein each draw is either a success or a failure. Moreover, the expected value of $X$ is $\mathbb{E}(X)=nN/K$.
@@
Set $V=V(G)$, and fix an arbitrary vertex $x\in V$. Let $S\subseteq V$ be drawn uniformly among all $s$-sets. We say that vertex $v\in V$ is bad for $S$ if $\overline{\deg}_{G[S\cup\{v\}]}(v)<\overline{\deg}_{G}(v)-\varepsilon$. Moreover, $S$ is good if it contains no vertex which is bad for $S$. So $P$ is the $s$-graph of good edges. It suffices to show that, conditioning on the event that $x\in S$, every vertex $v\in V$ is bad for $S$ with probability at most $\exp(-\Omega(s))$. Indeed, in this case,
@@
To obtain the above bound, consider a vertex $v\in V$ with $\delta=\overline{\deg}_{G}(v)-\varepsilon$. We can assume that $\delta\geq 0$ as otherwise there is nothing to show. Let $q$ be $1$ if $xv\in E(G)$ and 00 otherwise. Let $X=\deg_{G[S\cup\{v\}]}(v)$ conditioning on $x\in S$, and note that $X\sim q+\text{H}(n-1,\deg_{G}(v)-q,s-1)$. So $\mathbb{E}(X)\geq(s-1)(\deg_{G}(v)-q)/(n-1)\geq(\delta+(3/4)\varepsilon)(s-1)$. It follows by Lemma A.2 applied with $s-1$, $(\varepsilon/4)(s-1)$ playing the rôles of $n$, $\ell$ that
@@ B. Almost perfect tilings
In this section, we show Lemma 5.1 following the exposition of our work on hypergraphs [7, §7].
@@
We also require the following (crude) Dirac-type bound for hypergraph matchings due to Daykin and Häggkvist [2], whose short proof is included for sake of completeness.
@@
Consider a largest matching $M$ in $P^{\prime}$. We can assume that $M$ covers all but $s$ vertices of $P^{\prime}$. Otherwise just add some dummy edges to $M$. Let $v_{1},\dots,v_{s}$ be the vertices not covered by $M$, where $v_{i}$ is in cluster $V_{i}$ of $P^{\prime}$. Let $\sigma_{1},\dots,\sigma_{s}\in\mathcal{S}_{s}$ be arbitrary permutations with $\sigma_{i}(i)=i$. Let $X=(e_{1},\dots,e_{s-1})\in M^{s-1}$. For $\sigma_{i}$ and $1\leq j\leq s$ with $j\neq i$, let by $v_{\sigma_{i}}(j)$ be the vertex of $e_{\sigma(i)}$ in cluster $V_{j}$. Set $v_{\sigma_{i}}(i)=v_{i}$. We say that $X$ is good for index $i$, if $\{v_{\sigma_{i}}(1),\dots,v_{\sigma_{i}}(s)\}\in E(P^{\prime}).$ There are at most $m^{s-1}$ choices for $X$. Note that at most $\binom{s-1}{2}n^{s-2}<\varepsilon/2$ of theses choices contain a repeated matching edge. Moreover, for fixed $i$, there are at least $(1-1/s+\varepsilon/2)m^{s-1}$ good $X\in M^{s-1}$. (One for each $M$-partite neighbour of $v_{i}$.) So there is some $X$ without repeated matching edge, which is good for all $1\leq i\leq s$. Now we just have to choose $\pi_{1},\dots,\pi_{s}$ such that we may find a perfect matching between the vertices of $X$ and and $v_{1},\dots,v_{s}$. But this is an easy exercise. So we have a new matching $M^{\prime}$. If $M$ did not have dummy edges, we are done. Otherwise we have found a larger matching than $M$, which is a contradiction. ∎
@@
Let us briefly explain how we intend to prove Lemma B.3. Our plan is to extract additional regular tuples from the $(1-\lambda)n\geq\eta n$ vertices not covered by $Q_{1}$. We shall do this by using the assumption that the minimum degree is $\mu\binom{n-1}{s-1}$ above the threshold for a perfect matching. To track these gains, we define $\nu=\mu/16$ and $\beta=\nu\eta/2$. In a first step, we find a tiling $\mathcal{Q}_{\text{fresh}}$ that covers $4\beta\eta n$ additional vertices outside of $Q_{1}$ and some vertices inside of $Q_{1}$. In a second step, we ‘recycle’ the remainder of $Q_{1}$ using with another tiling $\mathcal{Q}_{\text{rec}}$ such that $\mathcal{Q}_{\text{fresh}}\cup\mathcal{Q}_{\text{rec}}$ together cover all but $2\beta\eta n$ vertices in $Q_{1}$. We note that the leftover comes from the fact that removing $\mathcal{Q}_{\text{fresh}}$ leaves parts of $Q_{1}$ a bit unbalanced. In any case, the relative difference $162\beta n=\mu\eta/$ 16 presents the share of additional covered vertices at the end of the argument.
@@
We set $\nu=\mu/16$ and $\beta=\nu\eta/2$. Moreover, let $m_{1}=m$, $m_{2}=(\mu/8)m_{1}$ and $m_{3}=m^{\prime}=\gamma 2\beta m_{2}$. We begin by defining family $\mathcal{U}$ of disjoint sets in $V(P)$, as follows. For each $s$-tuple in $Q_{1}$ with parts $\{V_{1},\dotsc,V_{s}\}$, pick a maximal number of vertex-disjoint sets of size $m_{2}$ inside each $V_{i}$, and add all of them to $\mathcal{U}$. It follows the sets of $\mathcal{U}$ cover all but $m_{2}=(\mu/8)m_{1}\leq(\mu/8)|V_{i}|$ vertices in each $V_{i}$. Next, we extend $\mathcal{U}$ by subdividing the vertices outside of $Q_{1}$ into a maximal number of vertex-disjoint $m_{2}$-sets such that $r=|\mathcal{U}|$ is divisible by $s$. This implies that outside of $V(Q_{1})$, the family $\mathcal{U}$ covers all but at most $sm_{2}$ vertices. We remark that $1/s,\mu$ are much larger than $1/r$ by the choice of $\alpha$. Now, let $R$ be an $s$-graph with vertex set $\mathcal{U}$ and an edge $X$ if $P$ contains at least $4\nu m_{2}^{s}$ $X$-partite edges. So $R$ plays the rôle of a reduced graph in the context of a Regularity Lemma. It is therefore not surprising that $R$ inherits the minimum degree condition of $P$.
@@
Since $r$ is divisible by $s$, the claim and Lemma B.2 imply that there is a perfect matching $\mathcal{M}$ in $R$. We will construct a $\mathcal{Q}_{2}$-tiling by finding tuples inside each edge $X\in\mathcal{M}$, as follows. Each edge $X$ of $\mathcal{M}$ corresponds to an $s$-partite family $\{U_{1},\dotsc,U_{s}\}$. We construct a $\mathcal{Q}_{2}$-tiling in $X$ greedily, by finding one tuple after another. Note that as long as the number of used vertices in each part $U_{i}$ is at most $3\nu m_{2}$, there remain at least $\nu m_{2}^{s}\geq dm_{2}^{s}$ edges among the leftover vertices. (Here we used that $d\leq\mu/16=\nu$.) Within the unused vertices in $X$, select subsets $U^{\prime}_{i}\subseteq U_{i}$, for all $1\leq i\leq s$, of size $\beta m_{2}$ each such that $d_{P}(U^{\prime}_{1},\dotsc,U^{\prime}_{s})\geq d$. (This can be done using an averaging argument). We apply Lemma B.1 with $\beta m_{2}$ playing the rôle of $m$, to obtain an $(\varepsilon,d)$-lower-regular $\tilde{m}$-balanced $s$-tuple, where $m_{3}=2\gamma\beta m_{2}\leq\tilde{m}\leq\beta m_{2}$. Note that this lower-regular $s$-tuple belongs to $\mathcal{Q}_{2}$. We iterate this, to greedily find a $\mathcal{Q}_{2}$-tiling on at least $3\nu sm_{2}$ vertices, and at most $3\nu sm_{2}+\beta m_{2}$ vertices, in the $X$-partite subgraph of $P$ hosted by $\bigcup X$.
@@
Next, we define a $\mathcal{Q}_{2}$-tiling $\mathcal{Q}_{\text{rec}}\subseteq P$ in $V(Q_{1})\smallsetminus V(\mathcal{Q}_{\text{fresh}})$ to ‘recycle’ what is left of $Q_{1}$. Let $\{V_{1},\dotsc,V_{s}\}$ be the parts of an ${\tilde{m}}$-balanced $(\varepsilon,d)$-lower-regular $s$-tuple in $Q_{1}$. Recall that, $\mathcal{Q}_{\text{fresh}}$ covers between $3\nu m_{2}$ and $(3\nu+\beta)m_{2}$ vertices in each part of $\mathcal{U}$. This implies in particular that the difference of leftover vertices between any two clusters of $V_{1},\dots,V_{s}$ is at most $\beta{\tilde{m}}$. We may therefore pick subsets $U_{1}\subseteq V_{1},\dotsc,U_{s}\subseteq V_{s}$ of size $\eta{\tilde{m}}$ of size $m^{\ast}=(1-3\nu-\beta){\tilde{m}}$ among the uncovered vertices. Note that $(U_{1},\dots,U_{s})$ is still $(2\varepsilon,d^{\prime})$-lower-regular for $d^{\prime}\geq d-\varepsilon\geq d/2$ and thus in $\mathcal{Q}_{2}$. Indeed, for all choices of $X_{1}\subseteq U_{1},$ $\dots$,… , $X_{s}\subseteq U_{s}$ of size at least $2\varepsilon m^{\ast}\geq\tilde{m}$ each, we have $d_{P}(X_{1},\dots,X_{s})\geq d-\varepsilon$ by $(\varepsilon,d)$-regularity of $(V_{1},\dots,V_{s})$. Let $\mathcal{Q}_{\text{rec}}\subseteq P$ be the $\mathcal{Q}_{2}$-tiling obtained this way. It follows that $\mathcal{Q}_{\text{rec}}\cup\mathcal{Q}_{\text{fresh}}$ covers all but $\beta v(Q_{1})=\beta\lambda n$ vertices of $V(Q_{1})$. It follows that $\mathcal{Q}_{\text{fresh}}\cup\mathcal{Q}_{\text{rec}}$ covers at least $(1-\beta)\lambda n+4\beta n\geq(\lambda+4\beta-2\beta)n=(\lambda+\nu\eta)n$ vertices. ∎

### v2

@@ 6. Connecting and rooting blow-ups
In this section, we show Lemmata 2.2 and 4.2, starting with the former. The approach follows the (classic) solution by Kővári, Sós and Turán [9] to Zarankiewicz’s problem, which is based on a double-counting argument.
@@
Note that every vertex in $U\cup V$ has at least $(1/2+\varepsilon/2)n$ neighbours in $W$. Let $W_{U}\subseteq W$ be the set of vertices of $W$ with at least $\varepsilon m/8$ neighbours in $U$. Counting edges between $U$ and $W$, we get $(1/2+\varepsilon/2)mn\leq|W_{U}|m+(n-|W_{U}|)\varepsilon m/8$, which gives
@@
Now let $W_{V}\subseteq W$ be the set of vertices of $W$ with at least $\varepsilon m/8$ neighbours in $V$. An analogous argument gives that $|W_{V}|\geq\left(1/2+\varepsilon/2\right)n$. Hence, the set $W^{\ast}=W_{U}\cap W_{V}$ has size at least $\varepsilon n$.
Set $t=m^{\prime}$. We claim that there are $t$-sets $W_{1}\subseteq U$, $W_{2}\subseteq V$ and $W_{3}\subseteq W$ such that $G$ contains the complete bipartite graphs $K(W_{1},W_{3})$ and $K(W_{2},W_{3})$. Suppose otherwise. Then no two $t$-sets in $U$ and $V$ have more than t1t-1 common neighbours in $W^{\ast}$. Let $\deg(w,U)$ denote the number of edges between $w$ and $U$. Therefore
@@
It remains to show how to obtain rooted blow-ups. The proof combines the ideas of Lemma 4.1 and the solutions to Zarankiewicz’s problem.
@@
Let $G^{\prime}$ be the graph obtained from $G$ by removing the edges inside $X$. Since $|X|=m_{1}\leq\varepsilon n/4$ this barely affects the minimum degree. So we still have $\delta(G^{\prime})\geq(1/2+3\varepsilon/4)n$. Let $P$ be the $s$-uniform hypergraph where an $s$-set forms an edge if it has exactly one vertex in $X$ and the induced subgraph $G^{\prime}[S]$ has minimum degree at least $(1/2+\varepsilon/2)s$. By Lemma 3.1 and crude bounding, we have $e(P)\geq|X|n^{s-1}/2$. Consider a random partition of $V(G)\smallsetminus X$ by assigning each vertex uniformly to one of s1s-1 parts. Together with $X$, this gives a partition of $V(G)$ into $s$ parts. Let $P^{\prime}\subseteq P$ be the $s$-partite $s$-graph, which contains the crossing edges. By a first moment argument, it follows that with positive probability $e(P^{\prime})\geq|X|n^{s-1}k!/(2k^{k})$. Fix such a partition for the rest of the proof. Recall that there are at most $2^{s^{2}}$ labelled graphs on $s$ vertices. Let $R$ be the labelled graph that appears on the edges of $P^{\prime}$ most frequently, where the labels correspond to the parts. We pass to a subgraph $P^{\prime\prime}\subseteq P^{\prime}$ keeping only the $s$-sets $S$ where $G[S]$ is isomorphic to $R$. So $P^{\prime\prime}$ has at least $|X|n^{s-1}k!/(k^{k}2^{s^{2}+2})$ edges. Note that by our choice we can assume that for each $S\in P^{\prime\prime}$, $G[S]$ is isomorphic to $R$, $|S\cap X|=1$ and also $G[S\smallsetminus X]$ is isomorphic to the same (s1)(s-1)-graph $R^{\prime}$, no matter the choice of $S$.
@@
Now let $1/s\gg\gamma_{1}\gg\gamma_{2}\gg c$, so that $e(P^{\prime\prime})\geq\gamma_{1}|X|n^{s-1}$. Let $\mathcal{R}^{\prime}$ be the set of ordered (s1)(s-1)-tuples in GXG-X, inducing a copy of $R^{\prime}$, which can be extended to at least $\gamma_{1}|X|/2$ copies of $R$ by adding a vertex of $X$. Since $e(P^{\prime\prime})\geq\gamma_{1}|X|n^{s-1}$, we have that $|\mathcal{R}^{\prime}|\geq\gamma_{1}n^{s-1}/2$. By Lemma 3.2, there is a blow-up $K\subseteq G-X$ of $R^{\prime}$ with all clusters of size $t=\gamma_{2}\log n$. Let $R^{\prime}_{1},\dotsc,R^{\prime}_{t}$ be vertex-disjoint copies of $R^{\prime}$ in the blow-up $K$. Form an auxiliary bipartite graph $B$ with vertex classes $X$ and $W=\{R^{\prime}_{1},\dotsc,R^{\prime}_{t}\}$ where $u\sim R^{\prime}_{i}$ if $R^{\prime}_{i}$ can be extended to a copy of $R$ by adding $u$. Then $B$ has at least $\gamma_{1}t|X|/2$ edges.
@@
To conclude, it is enough to find a complete bipartite graph in $B$ with parts of size $c\ln n$. The existence of such a graph follows from the usual bounds for Zarankiewicz’s problem. To be precise, let $p=c\log n$ and suppose $B$ does not have a complete bipartite graph with two classes of size $p$. Then each $p$-set $Y\subseteq W$ has at most p1p-1 common neighbours $N_{B}(Y)$ in $X$. Thus, double-counting and convexity reveal
$(p-1)\binom{t}{p}$ $\geq\sum_{Y\in\binom{W}{p}}|N_{B}(Y)|=\sum_{v\in X}\binom{\deg_{B}(v)}{p}$
$\geq|X|\binom{e(B)/|X|}{p}\geq|X|\binom{\gamma_{1}t/2}{p}\geq(\gamma_{1}/4)^{p}|X|\binom{t}{p}$,
@@
$p\geq(\gamma_{1}/4)^{p}|X|\geq|X|n^{-1/4}\geq n^{1/4},$
where the second inequality follows from the (small enough) choice of $c$, and the last inequality follows from $|X|=m_{1}\geq n^{1/2}$. This contradicts the choice of $p$, and proves the result. ∎
@@ Acknowledgements
We are grateful for the insightful comments of an anonymous referee, which helped to improve the presentation and correctness of the arguments.
Richard Lang was supported by the Ramón y Cajal programme (RYC2022-038372-I) and by grant PID2023-147202NB-I00 funded by MICIU/AEI/10.13039/501100011033. Nicolás Sanhueza-Matamala was supported by ANID-FONDECYT Iniciación Nº11220269 grant and ANID-FONDECYT Regular Nº1251121 grant.
@@ A. Inheritance
For a $k$-graph $G$ on $n$ vertices, we denote the relative degree of a vertex $v\in V(G)$ by $\overline{\deg}_{G}(v)=\deg_{G}(v)/(n-1)$. The following instance of the inheritance principle immediately implies Lemma 3.1.
@@
We show Lemma A.1 using the following standard tail bound. Recall that a hypergeometrically distributed random variable $X\sim\text{H}(N,K,n)$ describes the probability of $k$ successes in $n$ draws, without replacement, from a finite population of size $N$ that contains exactly $K$ objects with that feature, where each draw is either a success or a failure. Moreover, the expected value of $X$ is $\mathbb{E}(X)=nN/K$.
@@
Set $V=V(G)$, and fix an arbitrary vertex $x\in V$. Let $S\subseteq V$ be drawn uniformly among all $s$-sets. We say that a vertex $v\in V$ is bad for $S$ if $\overline{\deg}_{G[S\cup\{v\}]}(v)<\overline{\deg}_{G}(v)-\varepsilon$. Moreover, $S$ is good if it contains no vertex which is bad for $S$. So $P$ is the $s$-graph of good edges. It suffices to show that, conditioning on the event that $x\in S$, every vertex $v\in V$ is bad for $S$ with probability at most $\exp(-\Omega(s))$. Indeed, in this case,
@@
To obtain the above bound, consider a vertex $v\in V$ with $\delta=\overline{\deg}_{G}(v)-\varepsilon$. We can assume that $\delta\geq 0$ as otherwise there is nothing to show. Let $q$ be $1$ if $xv\in E(G)$ and 0 otherwise. Let $X=\deg_{G[S\cup\{v\}]}(v)$ conditioning on $x\in S$, and note that $X\sim q+\text{H}(n-1,\deg_{G}(v)-q,s-1)$. So $\mathbb{E}(X)\geq(s-1)(\deg_{G}(v)-q)/(n-1)\geq(\delta+(3/4)\varepsilon)(s-1)$. It follows by Lemma A.2 applied with s1s-1, $(\varepsilon/4)(s-1)$ playing the rôles of $n$, $\ell$ that
@@ B. Almost perfect tilings
In this section, we show Lemma 5.1 following the exposition of our work on hypergraphs [11, §7].
@@
We also require the following (crude) Dirac-type bound for hypergraph matchings due to Daykin and Häggkvist [3], whose short proof is included for the sake of completeness.
@@
Consider a largest matching $M$ in $P^{\prime}$. We can assume that $M$ covers all but $s$ vertices of $P^{\prime}$. Otherwise, just add some dummy edges to $M$. Let $v_{1},\dots,v_{s}$ be the vertices not covered by $M$, where $v_{i}$ is in cluster $V_{i}$ of $P^{\prime}$. Let $\sigma_{1},\dots,\sigma_{s}\in\mathcal{S}_{s}$ be arbitrary permutations with $\sigma_{i}(i)=i$. Let $X=(e_{1},\dots,e_{s-1})\in M^{s-1}$. For $\sigma_{i}$ and $1\leq j\leq s$ with $j\neq i$, let $v_{\sigma_{i}}(j)$ be the vertex of $e_{\sigma(i)}$ in cluster $V_{j}$. Set $v_{\sigma_{i}}(i)=v_{i}$. We say that $X$ is good for index $i$ if $\{v_{\sigma_{i}}(1),\dots,v_{\sigma_{i}}(s)\}\in E(P^{\prime}).$ There are at most $m^{s-1}$ choices for $X$. Note that at most $\binom{s-1}{2}n^{s-2}<\varepsilon/2$ of these choices contain a repeated matching edge. Moreover, for fixed $i$, there are at least $(1-1/s+\varepsilon/2)m^{s-1}$ good $X\in M^{s-1}$. (One for each $M$-partite neighbour of $v_{i}$.) So there is some $X$ without repeated matching edge, which is good for all $1\leq i\leq s$. Now we just have to choose $\pi_{1},\dots,\pi_{s}$ such that we may find a perfect matching between the vertices of $X$ and $v_{1},\dots,v_{s}$. But this is an easy exercise. So we have a new matching $M^{\prime}$. If $M$ did not have dummy edges, we are done. Otherwise, we have found a larger matching than $M$, which is a contradiction. ∎
@@
Let us briefly explain how we intend to prove Lemma B.3. Our plan is to extract additional regular tuples from the $(1-\lambda)n\geq\eta n$ vertices not covered by $Q_{1}$. We shall do this by using the assumption that the minimum degree is $\mu\binom{n-1}{s-1}$ above the threshold for a perfect matching. To track these gains, we define $\nu=\mu/16$ and $\beta=\nu\eta/2$. In a first step, we find a tiling $\mathcal{Q}_{\text{fresh}}$ that covers $4\beta\eta n$ additional vertices outside of $Q_{1}$ and some vertices inside of $Q_{1}$. In a second step, we ‘recycle’ the remainder of $Q_{1}$ using another tiling $\mathcal{Q}_{\text{rec}}$ such that $\mathcal{Q}_{\text{fresh}}\cup\mathcal{Q}_{\text{rec}}$ together cover all but $2\beta\eta n$ vertices in $Q_{1}$. We note that the leftover comes from the fact that removing $\mathcal{Q}_{\text{fresh}}$ leaves parts of $Q_{1}$ a bit unbalanced. In any case, the relative difference $2\beta n=\mu\eta/16$ presents the share of additional covered vertices at the end of the argument.
@@
We set $\nu=\mu/16$ and $\beta=\nu\eta/2$. Moreover, let $m_{1}=m$, $m_{2}=(\mu/8)m_{1}$ and $m_{3}=m^{\prime}=\gamma 2\beta m_{2}$. We begin by defining a family $\mathcal{U}$ of disjoint sets in $V(P)$, as follows. For each $s$-tuple in $Q_{1}$ with parts $\{V_{1},\dotsc,V_{s}\}$, pick a maximal number of vertex-disjoint sets of size $m_{2}$ inside each $V_{i}$, and add all of them to $\mathcal{U}$. It follows that the sets of $\mathcal{U}$ cover all but $m_{2}=(\mu/8)m_{1}\leq(\mu/8)|V_{i}|$ vertices in each $V_{i}$. Next, we extend $\mathcal{U}$ by subdividing the vertices outside of $Q_{1}$ into a maximal number of vertex-disjoint $m_{2}$-sets such that $r=|\mathcal{U}|$ is divisible by $s$. This implies that outside of $V(Q_{1})$, the family $\mathcal{U}$ covers all but at most $sm_{2}$ vertices. We remark that $1/s,\mu$ are much larger than $1/r$ by the choice of $\alpha$. Now, let $R$ be an $s$-graph with vertex set $\mathcal{U}$ and an edge $X$ if $P$ contains at least $4\nu m_{2}^{s}$ $X$-partite edges. So $R$ plays the rôle of a reduced graph in the context of a Regularity Lemma. It is therefore not surprising that $R$ inherits the minimum degree condition of $P$.
@@
Since $r$ is divisible by $s$, the claim and Lemma B.2 imply that there is a perfect matching $\mathcal{M}$ in $R$. We will construct a $\mathcal{Q}_{2}$-tiling by finding tuples inside each edge $X\in\mathcal{M}$, as follows. Each edge $X$ of $\mathcal{M}$ corresponds to an $s$-partite family $\{U_{1},\dotsc,U_{s}\}$. We construct a $\mathcal{Q}_{2}$-tiling in $X$ greedily, by finding one tuple after another. Note that as long as the number of used vertices in each part $U_{i}$ is at most $3\nu m_{2}$, there remain at least $\nu m_{2}^{s}\geq dm_{2}^{s}$ edges among the leftover vertices. (Here we used that $d\leq\mu/16=\nu$.) Within the unused vertices in $X$, select subsets $U^{\prime}_{i}\subseteq U_{i}$, for all $1\leq i\leq s$, of size $\beta m_{2}$ each such that $d_{P}(U^{\prime}_{1},\dotsc,U^{\prime}_{s})\geq d$. (This can be done using an averaging argument.) We apply Lemma B.1 with $\beta m_{2}$ playing the rôle of $m$, to obtain an $(\varepsilon,d)$-lower-regular $\tilde{m}-$balanced $s$-tuple, where $m_{3}=2\gamma\beta m_{2}\leq\tilde{m}\leq\beta m_{2}$. Note that this lower-regular $s$-tuple belongs to $\mathcal{Q}_{2}$. We iterate this, to greedily find a $\mathcal{Q}_{2}$-tiling on at least $3\nu sm_{2}$ vertices, and at most $3\nu sm_{2}+\beta m_{2}$ vertices, in the $X$-partite subgraph of $P$ hosted by $\bigcup X$.
@@
Next, we define a $\mathcal{Q}_{2}$-tiling $\mathcal{Q}_{\text{rec}}\subseteq P$ in $V(Q_{1})\smallsetminus V(\mathcal{Q}_{\text{fresh}})$ to ‘recycle’ what is left of $Q_{1}$. Let $\{V_{1},\dotsc,V_{s}\}$ be the parts of an $\tilde{m}$}-balanced $(\varepsilon,d)$-lower-regular $s$-tuple in $Q_{1}$. Recall that, $\mathcal{Q}_{\text{fresh}}$ covers between $3\nu m_{2}$ and $(3\nu+\beta)m_{2}$ vertices in each part of $\mathcal{U}$. This implies, in particular, that the difference of leftover vertices between any two clusters of $V_{1},\dots,V_{s}$ is at most $\beta{\tilde{m}}$. We may therefore pick subsets $U_{1}\subseteq V_{1},\dotsc,U_{s}\subseteq V_{s}$ of size $m^{\ast}=(1-3\nu-\beta){\tilde{m}}$ among the uncovered vertices. Note that $(U_{1},\dots,U_{s})$ is still $(2\varepsilon,d^{\prime})$-lower-regular for $d^{\prime}\geq d-\varepsilon\geq d/2$ and thus in $\mathcal{Q}_{2}$. Indeed, for all choices of $X_{1}\subseteq U_{1},$ $\dots,$ $X_{s}\subseteq U_{s}$ of size at least $2\varepsilon m^{\ast}\geq\varepsilon\tilde{m}$ each, we have $d_{P}(X_{1},\dots,X_{s})\geq d-\varepsilon$ by $(\varepsilon,d)$-regularity of $(V_{1},\dots,V_{s})$. Let $\mathcal{Q}_{\text{rec}}\subseteq P$ be the $\mathcal{Q}_{2}$-tiling obtained this way. It follows that $\mathcal{Q}_{\text{rec}}\cup\mathcal{Q}_{\text{fresh}}$ covers all but $\beta v(Q_{1})=\beta\lambda n$ vertices of $V(Q_{1})$. It follows that $\mathcal{Q}_{\text{fresh}}\cup\mathcal{Q}_{\text{rec}}$ covers at least $(1-\beta)\lambda n+4\beta n\geq(\lambda+4\beta-2\beta)n=(\lambda+\nu\eta)n$ vertices. ∎